# Yi-Fang Wang, Chia-Ming Lee, Yi-Shiuan Chou

### v2

@@ 3.6 Limitations and Discussion
This study introduces a novel approach, GRACE, to address the challenge of DeepFake video detection in the presence of noisy face sequences. GRACE leverages feature entanglement with sparse constraints and a graph convolutional network with graph Laplacian regularization to effectively exploit the spatial-temporal correlations in face sequences while suppressing the impact of noise and distortions. The experimental results demonstrate the efficacy of GRACE in handling noisy face sequences and achieving state-of-the-art performance on several benchmark datasets.
However, it is essential to acknowledge the limitations of the current study and discuss potential future directions. One limitation is that while GRACE has shown strong performance on the evaluated datasets, its effectiveness on cross-dataset scenarios, where the training and testing data come from different sources, has not been extensively explored. The robustness of GRACE to domain shifts and variations in noise characteristics across different datasets requires further investigation. Nonetheless, it is worth emphasizing that GRACE represents the first dedicated effort to tackle the problem of noisy face sequences in DeepFake video detection, which has been largely overlooked in previous research. The proposed methodology and insights from this study lay a solid foundation for future work in this important direction.
@@ 4 Conclusions
In this work, we proposed a robust and generalized Graph-Regularized Attentive Convolutional Entanglement (GRACE) approach for DeepFake video detection, specifically designed to address the challenges posed by noisy and unreliable face sequences. The proposed GRACE framework leverages spatiotemporal feature entanglement, graph convolutional networks, and graph Laplacian regularization to effectively capture discriminative features while mitigating the impact of invalid facial images. Extensive experiments on benchmark datasets, including FF++ [1], Celeb-DF [3], and DFDC [2], demonstrate the superior performance of GRACE compared to state-of-the-art methods, especially under noisy face sequences. The robustness of GRACE is further validated through real-world simulations involving adversarial attacks on face detectors. The proposed GRACE represents a significant step forward in robust and generalized DeepFake video detection under challenging conditions, contributing to the development of more reliable multimedia forensics techniques in the era of deepfakes.

### v3

@@ 4.5 Limitations and Discussion
This study introduces a novel approach, GRACE, to address the challenge of DeepFake video detection in the presence of noisy face sequences. GRACE leverages feature entanglement with sparse constraints and a graph convolutional network with graph Laplacian smoothing prior regularization to effectively exploit the spatial-temporal correlations in face sequences while suppressing the impact of noise and distortions. The experimental results demonstrate the efficacy of GRACE in handling noisy face sequences and achieving state-of-the-art performance on several benchmark datasets.
However, it is essential to acknowledge the limitations of the current study and discuss potential future directions. One limitation is that while our work has shown strong performance on the evaluated datasets, its effectiveness on cross-dataset scenarios, where the training and testing data come from different sources, has not been extensively explored. The robustness of GRACE to domain shifts and variations in noise characteristics across different datasets requires further investigation. Nonetheless, it is worth emphasizing that GRACE represents the first dedicated effort to tackle the problem of noisy face sequences in DeepFake video detection, which has been largely overlooked in previous research. The proposed methodology and insights from this study lay a solid foundation for future work in this important direction.
@@ 5 Conclusions
In this work, we proposed a robust and generalized Graph-Regularized Attentive Convolutional Entanglement approach for DeepFake video detection, specifically designed to address the challenges posed by noisy and unreliable face sequences. The proposed framework leverages spatiotemporal feature entanglement, graph convolutional networks, and graph Laplacian smoothing prior regularization to effectively capture discriminative features while mitigating the impact of invalid facial images. Extensive experiments demonstrate the superior performance and robustness of GRACE compared to state-of-the-art methods, especially under noisy face sequences and adversarial attacks. It represents a significant step forward in robust and generalized DeepFake video detection under challenging conditions, contributing to the development of more reliable multimedia forensics techniques in the era of DeepFakes.
Appendix of GRACE: Graph-Regularized Attentive Convolutional Entanglement with Laplacian Smoothing for Robust DeepFake Video Detection
Theorem.1: Convergence of Graph Laplacian Regularized GCN
Let $\mathcal{L}(\mathbf{Z})$ be the loss function for a Graph Convolutional Network (GCN) with Graph Laplacian regularization, where $\mathbf{Z}\in\mathbb{R}^{n\times d}$ represents the node features, $n$ is the number of nodes, and $d$ is the feature dimension. Suppose the loss function $\mathcal{L}(\mathbf{Z})$ is convex and smooth in a local neighborhood of the optimal solution $\mathbf{Z}^{*}$. The GCN propagation rule with Graph Laplacian regularization is given by:
where $\hat{\mathbf{L}}=\hat{\mathbf{D}}^{-\frac{1}{2}}\hat{\mathbf{A}}\hat{\mathbf{D}}^{-\frac{1}{2}}$ is the normalized Graph Laplacian matrix, $\hat{\mathbf{A}}$ is the adjacency matrix with self-loops, $\hat{\mathbf{D}}$ is the corresponding degree matrix, $\mathbf{W}^{(l)}$ is the weight matrix at layer $l$, and $\sigma$ is a non-expansive activation function satisfying $|\sigma(\mathbf{x})-\sigma(\mathbf{y})|\leq L_{\sigma}|\mathbf{x}-\mathbf{y}|$ for any $\mathbf{x},\mathbf{y}$, with Lipschitz constant $L_{\sigma}\leq 1$. If the following conditions hold:
  1. 1.
    The spectral norm of the normalized Graph Laplacian matrix $|\hat{\mathbf{L}}|$ is bounded by $\lambda_{\max}(\hat{\mathbf{L}})\leq 2$. The eigenvalues of $\hat{\mathbf{L}}$ lie in the interval $[0,2]$, with the smallest eigenvalue $\lambda_{\min}(\hat{\mathbf{L}})=0$ if and only if the graph is connected.
  2. 2.
    The weight matrices $\mathbf{W}^{(l)}$ have bounded spectral norms, i.e., $|\mathbf{W}^{(l)}|\leq B_{W}$ for all $l$, and their variations are appropriately constrained [57]. This assumption ensures stable learning dynamics and prevents exploding gradients.
  3. 3.
    The composition of functions in each GCN layer, given by $f^{(l)}(\mathbf{Z})=\sigma(\hat{\mathbf{L}}\mathbf{Z}\mathbf{W}^{(l)})$, is Lipschitz continuous with constant $L_{f}=L_{\sigma}\cdot\lambda_{\max}(\hat{\mathbf{L}})\cdot B_{W}$.
  4. 4.
    The iterates $\mathbf{Z}^{(l)}$ stay within a local neighborhood of the optimal solution $\mathbf{Z}^{*}$ where the loss function is convex. This assumption is reasonable when the GCN is properly initialized and the learning rate is appropriately chosen.
Then, the iterative updates of the node features using the Graph Laplacian regularized GCN propagation rule will converge to the optimal solution $\mathbf{Z}^{*}$, i.e.,
$\lim_{l\rightarrow\infty}|\mathbf{Z}^{(l)}-\mathbf{Z}^{*}|=0$ (13)
Moreover, the convergence rate is linear, satisfying:
$|\mathbf{Z}^{(l+1)}-\mathbf{Z}^{*}|\leq\rho|\mathbf{Z}^{(l)}-\mathbf{Z}^{*}|,$ (14)
where $\rho=L_{f}<1$ is the contraction factor.
The proof follows from the convergence analysis of iterative algorithms for convex optimization problems. Consider two consecutive iterations $\mathbf{Z}^{(l)}$ and $\mathbf{Z}^{(l+1)}$:
$|\mathbf{Z}^{(l+1)}-\mathbf{Z}^{*}|$ $=|f^{(l)}(\mathbf{Z}^{(l)})-f^{(l)}(\mathbf{Z}^{*})|$ (15)
$\leq L_{f}|\mathbf{Z}^{(l)}-\mathbf{Z}^{*}|,$ (16)
where $L_{f}=L_{\sigma}\cdot\lambda_{\max}(\hat{\mathbf{L}})\cdot B_{W}$ is the Lipschitz constant of the composition of functions in each GCN layer. Under the assumptions that $\mathbf{Z}^{(l)}$ is sufficiently close to $\mathbf{Z}^{*}$ and the learning dynamics are stable, we have $L_{f}<1$, which implies:
$|\mathbf{Z}^{(l+1)}-\mathbf{Z}^{*}|\leq\rho|\mathbf{Z}^{(l)}-\mathbf{Z}^{*}|,$ (17)
where $\rho=L_{f}<1$ is the contraction factor. This result shows that the iterative sequence ${\mathbf{Z}^{(l)}}$ converges linearly to the optimal solution $\mathbf{Z}^{*}$ within its local neighborhood [58].
The boundedness of the spectral norm and the eigenvalue properties of the normalized Graph Laplacian matrix $\hat{\mathbf{L}}$ ensure stable propagation of node features and prevent over-smoothing [57, 59]. In the context of DeepFake detection, this implies that the GCN can effectively capture discriminative spatial-temporal patterns while mitigating noise and distortions. The bounded spectral norms and constrained variations of the weight matrices $\mathbf{W}^{(l)}$ promote stable learning dynamics and prevent exploding gradients [57]. This assumption is crucial for training deep GCNs on large-scale DeepFake datasets, where the model needs to learn robust and generalizable representations. The Lipschitz continuity of the composition of functions in each GCN layer, with a contraction factor $\rho<1$, guarantees the convergence of the iterative feature updates to the optimal solution. This property ensures that the Graph Laplacian regularized GCN can effectively learn discriminative node embeddings for DeepFake detection, even in the presence of challenging conditions such as partial occlusions, temporal inconsistencies, and diverse manipulation techniques. In summary, the convergence theorem for Graph Laplacian regularized GCNs, along with its assumptions and practical implications, provides a strong theoretical foundation for the proposed method in the context of DeepFake detection. The theorem justifies the effectiveness of incorporating Graph Laplacian regularization into GCNs for learning robust and discriminative spatial-temporal representations from noisy and challenging DeepFake videos.
Appendix.1: Hyperparameters Selection
TABLE VI: Performance evaluation of the proposed GRACE with different hyperparameter settings using FF++ [1]. $g_{\text{dim}}$ and $g_{n}$ are the embedding dimension and number of layers of GCN, respectively; $N$ is the frames extracted from the video; $n_{\text{out}}$ is the number of neurons of FC; $\alpha$ stands for weights of sparsity.
$m_{r}$ $N$ Acc. F1 AUC $g_{n}$ Acc. F1 AUC
0.8 12 0.922 0.875 0.971 12 0.856 0.711 0.950
0.8 20 0.948 0.918 0.978 4 0.948 0.918 0.982
0.8 16 0.944 0.916 0.983 8 0.944 0.916 0.983
0.7 12 0.958 0.936 0.983 12 0.896 0.816 0.964
0.7 20 0.954 0.930 0.986 4 0.952 0.926 0.983
0.7 16 0.960 0.938 0.985 8 0.960 0.938 0.985
- $\alpha$ Acc. F1 AUC $g_{\text{dim}}$ Acc. F1 AUC
0.8 $1e^{-7}$ 0.928 0.887 0.978 600 0.896 0.805 0.966
0.8 $1e^{-6}$ 0.942 0.910 0.974 200 0.934 0.886 0.981
0.8 $1e^{-5}$ 0.944 0.916 0.983 400 0.944 0.916 0.983
0.7 $1e^{-7}$ 0.954 0.930 0.981 600 0.938 0.896 0.984
0.7 $1e^{-6}$ 0.944 0.914 0.979 200 0.942 0.905 0.987
0.7 $1e^{-5}$ 0.960 0.938 0.985 400 0.960 0.938 0.985
- $n_{\text{out}}$ Acc. F1 AUC
0.8 1024 0.926 0.871 0.969
0.8 3072 0.924 0.876 0.939
0.8 2048 0.944 0.916 0.983
0.7 1024 0.944 0.908 0.984
0.7 3072 0.926 0.880 0.968
0.7 2048 0.960 0.938 0.985
To achieve optimal performance and robustness, we conducted a comprehensive ablation study to investigate the impact of various hyperparameters on the proposed GRACE method. This analysis provides valuable insights into the design choices and trade-offs involved in developing an effective DeepFake video detection system for real-world scenarios with noisy face sequences. Table VI presents the performance comparison of GRACE under different hyperparameter settings, evaluated on the challenging FF++ dataset [1].
1-1: Number of Extracted Frames ($N$)
The number of frames employed during the training and testing phases is a crucial aspect of GRACE. While using a larger number of frames might intuitively improve performance, it also significantly increases the computational complexity. To strike an optimal balance, we investigated the impact of varying the number of extracted frames. As shown in Table VI, using $N=8$ frames results in the lowest computational complexity but slightly compromises performance in terms of Accuracy, Macro F1-Score, and AUC. Conversely, increasing the number of frames to $N=20$ achieves state-of-the-art performance for most masking ratios during testing. Considering the trade-off between effectiveness and efficiency, we recommend using $N=16$ frames as the optimal setting for GRACE.
1-2: Number of GCN Layers ($g_{n}$)
The depth of the Graph Convolutional Network (GCN) plays a vital role in learning robust feature representations. However, stacking too many layers with the Graph Laplacian smooth prior may lead to over-smoothing of nodes and reduce the discriminative power. We explored the impact of varying the number of GCN layers ($g_{n}$) in GRACE. As presented in Table VI, setting $g_{n}=12$ results in suboptimal performance compared to $g_{n}=8$ and $g_{n}=4$, likely due to convergence difficulties within the given 200 epochs. While $g_{n}=4$ achieves outstanding performance overall, it slightly underperforms in highly noisy conditions (i.e., $m_{r}=0.8$) compared to $g_{n}=8$. Therefore, we suggest using $g_{n}=8$ as a balanced choice for stable and robust performance across various noise levels.
1-3: Sparsity Penalty Term ($\alpha$)
The sparsity penalty term $\alpha$ in the proposed loss function controls the balance between the sparsity constraint and the classification objective. A higher value of $\alpha$ encourages GRACE to learn a sparser feature representation, which is particularly beneficial for DeepFake video detection in the presence of invalid facial images. We investigated the impact of $\alpha$ by varying its value from $1e^{-7}$ to $1e^{-5}$. As shown in Table VI, a higher sparsity penalty enhances the network’s ability to learn essential and discriminative features, thereby reducing the influence of invalid faces and improving overall performance. However, setting $\alpha$ higher than $1e^{-5}$ leads to convergence difficulties. Based on our analysis, we recommend using $\alpha=1e^{-5}$ to achieve a balanced trade-off between sparsity and convergence stability.
1-4: GCN Embedding Dimension ($g_{\text{dim}}$)
The embedding dimension of the GCN ($g_{\text{dim}}$) determines the richness of the learned feature representations for DeepFake video detection. We investigated the impact of $g_{\text{dim}}$ by comparing the performance of GRACE with $g_{\text{dim}}\in{200,400,600}$, as shown in Table VI. Since the dimension of the graph representation ${\bm{A}}$ is $400\times 400$, intuitively, the best performance is achieved when $g_{\text{dim}}=400$. Reducing $g_{\text{dim}}$ below this value limits the expressive power of the GCN, while increasing it beyond introduces redundancy and harms performance. Therefore, we suggest setting $g_{\text{dim}}=400$ for optimal results.
1-5: Number of Fully Connected Layer Neurons ($n_{\text{out}}$)
To aggregate the output of the GCN and feed it into the softmax classifier, a simple fully connected (FC) layer is employed, projecting the graph representation to an $n_{\text{out}}$-dimensional feature vector. We investigated the impact of $n_{\text{out}}$ by comparing the performance of GRACE with $n_{\text{out}}\in{1024,2048,3072}$, as shown in Table VI. While $n_{\text{out}}=2048$ achieves excellent performance under highly noisy face sequences, the performance gap between $n_{\text{out}}=2048$ and $n_{\text{out}}=1024$ is insignificant, suggesting that the choice of $n_{\text{out}}$ is not highly sensitive. Based on our analysis, we recommend setting $n_{\text{out}}=2048$ for a good balance between performance and computational complexity.